# Human2LocoMan: Learning Versatile Quadrupedal Manipulation with Human Pretraining

### v1

@@ III-C Modularized Cross-embodiment Transformer
Although we employ a consistent data format and aligned input/output representations across embodiments, some modalities are not present or available for all embodiments. For example, the human operator is not equipped with a wrist camera, while the LocoMan robot has a wrist camera in some tasks to improve manipulation accuracy. In this case, we use masks defined during data collection to signify redundant dimensions in the observations as well as in the action labels. We refer the reader to Appendix Section VI-A for more implementation details.
@@ IV-A3 Data collection
For Toy Collection and Shoe Rack Organzation, we pretrain a model that utilizes the human data of both the unimanual and bimanual versions of the task, then we finetune the model on each unimanual or bimanual task with the corresponding robot data. For each task, we choose a set of training hyperparameters (e.g. batch size, chunk size) that are kept the same for all methods. (See Appendix Section VI-C.) We also list the model hyperparameters we use for our method and the baselines in the Appendix Section VI-A and VI-B.
@@ IV-A5 Baselines
More implementation details of these baselines can be found in Appendix Section VI-B. For the HPT baseline, we train with several different settings: training with only LocoMan data, pretraining with our human data and finetuning on LocoMan data, and directly finetuning the released HPT checkpoints with LocoMan data. For the HIT baseline, we only train on LocoMan data, as it is unable to incorporate human data.
@@ IV-B Results and Analysis
Task performance. We summarize the success rates and task scores of our method and HIT across all tasks in Table II. Human2LocoMan’s MXT achieves strong performance on all tasks using a relatively small dataset. The baseline method also attains decent performance on most tasks. These results highlight the high quality of our collected data and demonstrate the effectiveness of Human2LocoMan’s data collection and training pipeline.
@@
Compared to HIT. As shown in Table II, in most evaluated tasks, spanning both unimanual and bimanual modes and across both ID and OOD inference scenarios, MXT without pretraining achieves comparable or superior performance relative to HIT. Moreover, pretrained MXT consistently outperforms the HIT baseline in terms of both success rate and task score. From Figure6, we find that MXT demonstrates lower validation loss compared to HIT on most tasks, indicating superior training convergence. The performance improvement is particularly evident in tasks with larger datasets, suggesting that MXT scales more effectively with increasing data availability. Notably, HIT achieves a significantly lower validation loss compared to the MXT variants, while attaining comparable performance in SR and TS metrics under both ID and OOD settings relative to the best MXT model. As shown in the substep success analysis in Figure 5, the primary failures of the lower-performing MXT models occur during the first two substeps, push” and tap1.” One potential reason for this is that the unimanual shoe organization task exhibits relatively less variation in object locations and types compared to other tasks, which may favor HIT despite its lack of modular designs and pretraining.
Compared to HPT. From Figure 4, we observe that MXT consistently outperforms HPT in both SR and TS metrics across all combinations of pretraining and data sizes on the toy collection tasks. Validation loss results, shown in Figure 7, reveal a similar trend in the unimanual toy collection task across a broader range of dataset sizes. Notably, we observe severe overfitting in HPT experiments when training on our datasets, a phenomenon not observed in MXT. This further suggests that the modular design of the MXT architecture facilitates better generalization.
@@
Efficiency, robustness, and generalizability. As shown in Table II, pretraining on human data has a substantial positive impact on LocoMan manipulation performance. The policy maintains strong performance even when robot data is limited, highlighting both its efficiency and robustness. We hypothesize that MXT benefits from learning useful complementarities—i.e., positive transfer effects—between human demonstrations and LocoMan robot data. Specifically, comparing MXT-Pretrained to MXT-Scratch in Table II, we observe that pretraining improves performance on TC-Uni, TC-Bi, and Scooping tasks under ID settings, where objects exhibit diverse locations. MXT-Pretrained tends to produce smoother and more robust motions, enabling more accurate localization of target objects. For instance, as shown in Figure 5, MXT-Pretrained achieves substantially better scooping performance—which requires precise localization—compared to all other methods. Moreover, Table II reveals large performance gaps on OOD objects in tasks such as TC-Bi, SO-Uni, and Pouring, where OOD objects differ significantly from ID objects in shape, texture, and color. These results suggest that MXT, by leveraging human demonstrations during the pretraining stage, is able to generalize effectively to novel scenarios unseen during robot training.
@@
Our framework presents positive cross-embodiment transfer despite substantial embodiment gaps. From Figure 7, we see the gap in validation loss between HPT-Pretrained and HPT-Scratch is not as much as for MXT. The HPT-Small and HPT-Base models also do not generalize as well as MXT-Pretrained. This highlights the ability of MXT to consume human data which has a large embodiment gap from the LocoMan data.
For more concrete comparisons, we present SR and TS results based on 36 trials, comprising 24 OOD and 12 ID trials, as shown in Figure 4. HPT consistently underperforms compared to MXT, both when finetuned and when trained from scratch. We attribute this performance gap to a combination of HPT’s lack of modular design and its use of frozen image encoders. To further investigate the impact of modularity, we introduce an ablation variant of MXT, referred to as MXT-Agg, in which we aggregate the input modalities: a single visual tokenizer is used to encode all visual observations, a single proprioceptive tokenizer for all proprioceptive inputs, and a single detokenizer for all action dimensions—mirroring HPT’s design. MXT-Agg incorporates HPT’s key features, including cross-attention tokenization and trunk weight sharing, while still finetuning the vision encoders and remaining architecturally comparable to MXT. Across evaluations, MXT consistently benefits from pretraining and outperforms MXT-Agg when both are finetuned, highlighting the advantages of modularized tokenization for effectively leveraging human data.
Notably, MXT-Agg sometimes transfers suboptimally with respect to HPT, as evidenced by little to no improvement when finetuning the pretrained model compared with training from scratch. This is likely due to increased representation power in the tokenizer, which permits more overfitting in the transformer trunk and could negatively impact the trunk transferability. However, with the incorporation of our modular design, MXT is trained with additional regularization and exhibits improved transferability. The modular design effectively aids in the trade-off between more network representation power and better transferability in our framework, and allows attaining both qualities.
@@ V Limitations
While our system introduces a novel approach to cross-embodiment manipulation and efficient data collection for quadrupedal robots, it has several limitations. First, the teleoperation system still requires some practice for human operators to achieve precise manipulations and may feel unintuitive in certain aspects, such as controlling torso movements via head motions. Second, although we envision our system enabling large-scale cross-embodiment learning, in this work we have not yet scaled it to other robotic platforms or incorporated additional robotic datasets. As future work, we plan to validate its scalability and robustness across different robot types, including robotic arms and humanoids.
@@ VI Conclusion
In this paper, we introduce Human2LocoMan, a unified framework for flexible data collection and cross-embodiment learning to enable versatile quadrupedal manipulation skills on the open-source LocoMan platform. Our teleoperation and human data collection systems allow efficient acquisition of large-scale, high-quality datasets by bridging the action spaces between human and robot embodiments. Built upon this foundation, we propose Modularized Cross-embodiment Transformer, a modular policy architecture that supports positive transfer from human demonstrations to robot policies. Through extensive experiments on six challenging household tasks, we demonstrate that Human2LocoMan enables strong performance, efficient training, and robust generalization to out-of-distribution scenarios, outperforming strong imitation learning baselines. Our results highlight the effectiveness of cross-embodiment learning and modularized policy design in advancing scalable, versatile quadrupedal manipulation.
APPENDIX
VI-A Implementation and Training details of MXT
Training Details. We list the training optimizer and the transformer trunk hyperparamters in Table IV. These hyperparameters are kept the same for all our experiments.
TABLE IV: MXT trunk and training hyperparameters
Hyperparameters Value
optimizer AdamW
learning rate 5e-5 (finetuning/from scratch)
1e-4 (pretraining)
scheduler constant
weight decay 1e-4
trunk encoder layers 4
trunk decoder layers 4
hidden dim 128
transformer feedforward dim 256
#attention heads 16
Cross Attention in Tokenizers and Detokenizers. In the tokenizers of MXT, we use a simple cross attention mechanism to transform the input feature of indefinite length into a fixed number of tokens. For the attention layer in all tokenizers, the hidden dim is 128, the number of attention heads is 4, each with a head dimension of 32, and the dropout rate is 0.1. Other hyperparameters of each tokenizer are shown in Table V.
Similarly, we also use cross attention to decode the action modalities in detokenizers from a fixed number of output transformer tokens. For the attention layer, the number of attention heads is 4, each with a head dimension of 16, and the dropout rate is 0.1. Other hyperparameters of each detokenizer are shown in Table VI
TABLE V: MXT tokenizer hyperparameters
Modality Input dimensions #tokens MLP widths
main image (3, 480 1280) 16 N/A
wrist image (3, 480, 640) 8
body pose (6,) 4 [128, 128]
EEF pose (12,) 4
EEF to body pose (12,) 4
gripper angles (2,) 4
TABLE VI: MXT detokenizer hyperparameters
Modalities Output dimensions #tokens
body pose (6,) 6
EEF pose (12,) 6
gripper angle (2,) 6
Masks for aligning embodiment modalities. We mentioned that masks are needed to exclude redundant dimensions or modalities that are not present in some embodiment, and here we give a more detailed description of our implemented masks.
a) Masks on images. We recognize that some image view are not available for all embodiments and tasks. In our current framework, we assume there are at most two camera views (or image modalities): the main camera and the wrist camera. However, this can be easily extended within our framework to cater to any number of camera views. When one of these camera views is not present, we directly mark this in the transformer mask of the trunk and fill in dummy tokens in the corresponding positions, so that the positions associated with this image modality will not be attended on.
b) Masks on proprioceptive states. In some cases, the proprioceptive states may have some or all dimensions that should not be considered for the task. For example, in single-arm tasks, the poses of the left end effector, or the last half of the end effector pose modality, will not be considered, and in bimanual tasks where the LocoMan body is upright, the body pose is fixed and therefore redundant in the observations. When part of a proprioception modality are redundant dimensions, we apply zero padding on these dimension and perform encoding through the tokenizer as usual. Different from how we treated masked image modalities, this has no effect on the transformer mask of the trunk. When an entire proprioception modality should be disregarded, however, we handle this modality in a similar to the image modalities and apply the transformer mask accordingly.
Data Normalization. For both human and LocoMan data, we apply data normalization on observations and action labels. For non-image data, we estimate the per-dimension mean the standard deviation from the dataset, and normalize the data with the usual approach:
$\bar{x}_{t}=\frac{x_{t}-\text{mean}}{\text{std}}.$
For image data, the mean and standard deviation are set as the ImageNet statistics for the RGB channels: mean = [0.485, 0.456, 0.406], and std = [0.229, 0.224, 0.225]. The images are normalized in the same way with these parameters.
Dropout in Pretraining. We discover that increasing the dropout in transformer trunk can improve the finetuning performance for MXT. In practice, we find that setting the pretraining dropout rate to 0.5 for scooping and 0.4 for all other tasks yield reasonably good performance. When training with LocoMan data, including training from scratch and finetuning, the transformer trunk dropout rate is reverted to 0.1.
TABLE VII: HIT hyperparameters
Hyperparameters Value
optimizer AdamW
learning rate 2e-5
scheduler constant
weight decay 1e-4
encoder layers 4
decoder layers 4
hidden dim 128
#attention heads 8
feature loss weight 0.001
image backbone ResNet18
TABLE VIII: HPT hyperparameters
Hyperparameters Value
optimizer AdamW
learning rate 5e-5 (finetuning/from scratch)
1e-4 (pretraining)
scheduler constant
weight decay 1e-4
trunk
#transformer blocks 16
hidden dim 128
feedforward dim 256
#attention heads 8
action head
#attention heads 8
head dim 64
dropout 0.1
output dim 20
image stem
encoder ResNet18
MLP widths [128]
#tokens 16
state stem
MLP widths [128]
#tokens 16
VI-B Implementation details of baselines
HIT. Our implementation of Humanoid Imitation Transformer [20] is based on the released codebase, with only minor modifications to accommodate our data format. The hyperparameters used for training are summarized in Table VII.
HPT. We follow the original implementation of HPT [78], with the main exception that we changed the data normalization method so as to align with the approach of other frameworks and to have a fair comparison of the validation loss. The hyperparameters we used when training HPT are summarized in Table VIII.
VI-C Global task-specific training parameters
We choose a set of training parameters for each specific task, and we keep these settings aligned across all methods as listed in Table IX.
TABLE IX: Global training parameters for each task
Task Mode Batch Size Training Steps Chunk Size
Toy Collection Unimanual 16 60000 60
Bimanual 16 60000 60
Shoe Organization Unimanual 24 80000 180
Bimanual 24 100000 120
Scooping Unimanual 24 100000 120
Pouring Bimanual 24 80000 180

### v2

@@ III-C Modularized Cross-embodiment Transformer
Although we employ a consistent data format and aligned input/output representations across embodiments, some modalities are not present or available for all embodiments. For example, the human operator is not equipped with a wrist camera, while the LocoMan robot has a wrist camera in some tasks to improve manipulation accuracy. In this case, we use masks defined during data collection to signify redundant dimensions in the observations as well as in the action labels. We refer the reader to Appendix Section -A for more implementation details.
@@ IV-A3 Data collection
For Toy Collection and Shoe Rack Organzation, we pretrain a model that utilizes the human data of both the unimanual and bimanual versions of the task, then we finetune the model on each unimanual or bimanual task with the corresponding robot data. For each task, we choose a set of training hyperparameters (e.g. batch size, chunk size) that are kept the same for all methods. (See Appendix Section -C.) We also list the model hyperparameters we use for our method and the baselines in the Appendix Section -A and -B.
@@ IV-A5 Baselines
More implementation details of these baselines can be found in Appendix Section -B. For the HPT baseline, we train with several different settings: training with only LocoMan data, pretraining with our human data and finetuning on LocoMan data, and directly finetuning the released HPT checkpoints with LocoMan data. For the HIT baseline, we only train on LocoMan data, as it is unable to incorporate human data.
@@ IV-B Results and Analysis
Task performance. We summarize the success rates and task scores of our method and HIT across all tasks in Table II. Human2LocoMan’s MXT presents strong performance using a relatively small robot dataset, achieving success rates above $79\%$ across all tasks. The baseline method also attains decent performance on most tasks. These results highlight the high quality of our collected data and demonstrate the effectiveness of Human2LocoMan’s data collection and training pipeline.
@@
Compared to HIT. As shown in Table II, in most evaluated tasks, spanning both unimanual and bimanual modes and across both ID and OOD inference scenarios, MXT without pretraining achieves comparable or superior performance relative to HIT. Moreover, pretrained MXT consistently outperforms the HIT baseline in terms of both SR and TS. Specifically, pretrained MXT achieves an average SR improvement of $41.9\%$ overall and $79.7\%$ under OOD settings, as well as an average TS improvement of $28.9\%$ overall and $51.4\%$ under OOD settings. From Figure 6, we find that MXT demonstrates lower validation loss compared to HIT on most tasks, indicating superior training convergence. The performance improvement is particularly evident in tasks with larger datasets, suggesting that MXT scales more effectively with increasing data availability. Notably, HIT achieves a significantly lower validation loss compared to the MXT variants in unimanual shoe organization, while attaining comparable performance in SR and TS metrics under both ID and OOD settings relative to the best MXT model. As shown in the substep success analysis in Figure 5, the primary failures of the lower-performing MXT models occur during the first two substeps, push” and tap1.” One potential reason for this is that the unimanual shoe organization task exhibits relatively less variation in object locations and types compared to other tasks, which may favor HIT despite its lack of modular designs and pretraining.
Compared to HPT. As shown in Figure 4, we present SR and TS results based on 36 trials, comprising 24 OOD and 12 ID trials. HPT consistently underperforms compared to MXT, both when finetuned and when trained from scratch, across all data sizes on the toy collection tasks. We attribute this performance gap to a combination of HPT’s lack of modular architecture and its use of frozen image encoders. Validation loss results, shown in Figure 7, reveal a similar trend in the unimanual toy collection task across a broader range of dataset sizes. Notably, HPT-Small and HPT-Base fail to achieve lower validation loss compared to HPT-Scratch, even pretrained on the large public dataset. Their policy rollout performance are evidently worse than HPT-Scratch and HPT-Pretrained so we did not scale up the real-robot evaluation. This indicates the challenge posed by the large embodiment gap and underscores the effectiveness of the human data collected by our system. Furthermore, we observe severe overfitting in HPT experiments when training on our datasets, a phenomenon not observed in MXT. This further suggests that the modular design of the MXT architecture facilitates better generalization.
@@
Efficiency, robustness, and generalizability. Summarizing from Table II, pretraining with human data improves the MXT policy SR by $38.6\%$ overall and $82.7\%$ under OOD settings, and boosts TS by $24.1\%$ overall and $58.9\%$ under OOD settings. This enables consistently stronger performance with only half the amount of robot data, demonstrating both the efficiency and robustness of our system. We hypothesize that MXT benefits from learning useful complementarities—i.e., positive transfer effects—between human demonstrations and LocoMan robot data. Specifically, comparing MXT-Pretrained to MXT-Scratch in Table II, we observe that pretraining improves performance on TC-Uni, TC-Bi, and Scooping tasks under ID settings, where objects exhibit diverse locations. MXT-Pretrained tends to produce smoother and more robust motions, enabling more accurate localization of target objects. For instance, as shown in Figure 5, MXT-Pretrained achieves substantially better scooping performance—which requires precise localization—compared to all other methods. Moreover, Table II reveals large performance gaps on OOD objects in tasks such as TC-Bi, SO-Uni, and Pouring, where OOD objects differ significantly from ID objects in shape, texture, and color. These results suggest that MXT, by leveraging human demonstrations during the pretraining stage, is able to generalize effectively to novel scenarios unseen during robot training.
@@
We have shown in Figure 4 that MXT outperforms HPT under both finetuning and training-from-scratch settings, benefiting from its modularized design and unfrozen image encoders in its architecture. Here, we observe the delta performance brought by human data across different methods. While both MXT and HPT benefit from pretraining on human data, MXT exhibits more effective Human-to-LocoMan transfer, achieving larger gains, especially in low-data and OOD settings. Additionally, as depicted in Figure 8, MXT-Pretrained consistently achieves lower validation loss than MXT-Scratch, whereas the gap between HPT-Pretrained and HPT-Scratch is less consistent and does not always show the same trend. These results highlight the ability of MXT to consume human data, despite the large embodiment gap with LocoMan.
To further investigate the impact of modularity, we introduce an ablation variant of MXT, referred to as MXT-Agg, in which we aggregate the input modalities: a single visual tokenizer is used to encode all visual observations, a single proprioceptive tokenizer for all proprioceptive inputs, and a single detokenizer for all action dimensions—mirroring HPT’s design. MXT-Agg incorporates HPT’s key features, including cross-attention tokenization and trunk weight sharing, while still finetuning the vision encoders and remaining architecturally comparable to MXT. Across evaluations, MXT consistently benefits from pretraining and outperforms MXT-Agg when both are finetuned, highlighting the advantages of modularized tokenization for effectively leveraging human data. Notably, MXT-Agg exhibits weaker transfer performance with respect to MXT and HPT, as evidenced by little to no improvement when finetuning the pretrained model compared with training from scratch. This is likely due to increased representation power in the tokenizer, which permits more overfitting in the transformer trunk and could negatively impact the trunk transferability. However, with the incorporation of our modular design, MXT is trained with additional regularization and exhibits improved transferability. The modular design effectively aids in the trade-off between more network representation power and better transferability in our framework, and allows attaining both qualities.
@@ V Limitations
While our system introduces a novel and effective approach for cross-embodiment learning and data collection in quadrupedal manipulation, it has several limitations. First, although the teleoperation interface enables whole-body control and flexible data collection, it still requires a learning curve for operators. In particular, controlling the robot’s torso via head motions may feel unintuitive and demand some practice to achieve high precision. Second, despite the modular design and cross-embodiment capability of our architecture, this work focuses solely on the LocoMan platform. We have not yet validated the generality of our system across a broader range of robot embodiments, such as table-top or mobile arms and humanoid robots. Lastly, while we have demonstrated effective human-to-robot transfer, we rely on a pretraining-finetuning pipeline for single robotic tasks, and future work could explore co-training with human data in multi-task settings. Extending Human2LocoMan to diverse robots, larger-scale datasets, and more learning paradigms remains promising future directions.
@@ VI Conclusion
We present Human2LocoMan, a unified framework for efficient data collection and cross-embodiment learning, enabling versatile quadrupedal manipulation skills on the open-source LocoMan platform. Our system integrates XR-based teleoperation with aligned human and robot data collection, supporting a shared observation-action representation across embodiments. To leverage this data effectively, we introduce the Modularized Cross-embodiment Transformer, a modular policy architecture designed for scalable pretraining and fine-tuning across different embodiments and manipulation modes. Extensive real-world experiments across six challenging household tasks demonstrate that Human2LocoMan achieves strong performance, robust generalization to out-of-distribution settings, and high data efficiency, attaining high success rates with only a small amount of robot data. Comparisons against strong baselines highlight the effectiveness of our modular design and the benefits of cross-embodiment pretraining using human data collected through our system.